# From Competition to Centralization: The Oligopoly in Ethereum Block Building Auctions

### v1

@@
Abstract.
The Ethereum block production process has evolved with the introduction of an auction-based mechanism known as Proposer-Builder Separation (PBS), allowing validators to outsource block production to builders and reap Maximal Extractable Value (MEV) revenue from builder bids in a decentralized market. In this market, builders compete in MEV-Boost auctions to have their blocks selected and earn potential MEV rewards. This paper employs empirical game-theoretic analysis to explore builders’ strategic bidding incentives in MEV-Boost auctions, focusing on how advantages in network latency and access to MEV opportunities affect builders’ bidding behaviors and auction outcomes. Our findings confirm an oligopolistic dynamic, where a few dominant builders, leveraging their advantages in latency and MEV access, benefit from an economy of scale that reinforces their market power, leading to increased centralization and reduced auction efficiency. Our analysis highlights the importance of fair MEV distribution among builders and the ongoing challenge of enhancing decentralization in the Ethereum block building market.
Key words and phrases:
Decentralized Finance, Ethereum Blockchain, Maximal Extractable Value, MEV-Boost Auction, Empirical Game-Theoretic Analysis
1. Introduction
Decentralized Finance (DeFi) has rapidly transformed the financial industry by leveraging blockchain technology to offer permissionless and decentralized services. Platforms like Ethereum Buterin et al. (2014) have been at the forefront of this revolution. Unlike traditional financial systems that rely on centralized authorities, DeFi operates transparently through a peer-to-peer network of participants, known as validators, who collectively process and record transactions into blocks on a distributed ledger.
This decentralized approach, while innovative, introduces unique challenges. Validators, motivated by self-interest, may exploit their roles to maximize personal profits in ways that undermine network decentralization and security. One such exploitation is infamously known as Maximal Extractable Value (MEV), which refers to the value validators can extract by manipulating transaction inclusion, exclusion, and ordering in the blocks they produce Daian et al. (2019). MEV can lead to substantial centralization risks, as validators with greater resources—both in capital and computational power—are better positioned to extract higher MEV, thereby gaining disproportionate influence over the network Ethereum (2024); Bahrani et al. (2024); Ethereum (2023).
To mitigate these risks and promote a fair distribution of MEV among validators, Ethereum introduced the Proposer-Builder Separation (PBS) mechanism Ethereum (2024). PBS decouples the block proposal and block construction processes, allowing the proposer—a validator selected by the protocol to add a new block—to outsource the task of block building and MEV extraction to specialized entities known as builders. In the current PBS implementation, the proposer acts as an auctioneer, auctioning off the block building task to builders through the an English-style auction, called MEV-Boost Flashbots (2022). Builders compete in an open market to produce the most valuable blocks, submitting their proposals along with bids to intermediaries called relays. The proposer concludes the auction by selecting the block with the highest bid from the relays, thereby reaping a significant portion of MEV revenue through the auction. The winning builder also realizes MEV profits, with more than 90% blocks being currently built by builders via MEV-Boost Wahrstätter (2023).
During the auction process, builders’ bids are based on two main revenue sources: public transactions and private orderflow. Public transactions include user transactions accessible to all builders, offering tips (transaction fees) and MEV opportunities. Private orderflow consists of transactions and bundles sent directly to builders by providers like decentralized applications and MEV searchers, who detect MEV opportunities and create bundles to extract them. These private transactions, along with tips and payments [^1] https://docs.flashbots.net/flashbots-auction/advanced/coinbase-payment, are delivered either exclusively or to multiple builders for block inclusion. Builders then assemble public and private transactions into blocks and submit bids based on block value.
In an ideal market, builders have equal access to orderflow and network resources, creating a competitive, decentralized block-building ecosystem where MEV is efficiently captured via the MEV-Boost auction. However, in reality, builders seek a competitive edge by optimizing their connections with orderflow providers and relays, focusing on reducing latency and securing greater orderflow access. Some builders gain exclusive access to private transactions, and larger builders often vertically integrate by incorporating their own MEV searchers and relays. While this strengthens their bidding power, vertical integration has also led to increasing centralization, with two dominant builders (beaverbuild and Titan) winning 95% of auctions, resulting in an oligopoly Wahrstätter (2023).
Extensive literature has explored the impact of vertical integration on builders’ performance and market centralization Always (2023); Gupta et al. (2023); Kilbourn (2022, 2023); Yang et al. (2024). However, practical insights into how advantages in latency and access to private orderflow influence builders’ strategic bidding behaviors in MEV-Boost auctions remain limited. A previous study Wu et al. (2024b) demonstrated that under idealized market conditions, builders are incentivized to engage in a form of collusion, marginally outbidding each other in a way that leads to price-fixing within the auction. Nonetheless, the strategy space in that analysis was significantly constrained, relying on only three exemplary bidding strategies—truthful, adaptive, and last-minute—raising questions about the robustness of the results.
In this paper, we address these questions by conducting an empirical game-theoretic analysis of the MEV-Boost auction using the MEV-Boost auction model Wu et al. (2024a) with a richer strategy space. We employ a generalized empirical method known as the meta game approach, which involves constructing a higher-level game where strategies are combinations or distributions over base-level strategies, termed meta strategies, allowing us to capture a more comprehensive range of builders’ strategic bidding behaviors (Section 3). Utilizing meta strategies, we investigate builders’ strategic bidding incentives under various scenarios (Section 4) and analyze how advantages in latency and access to private orderflow impact their incentives and the overall auction outcomes (Section 5). We finally discuss the challenges in the Ethereum block building market based on our analyses and limitations of our approach (Section 6). The paper makes the following contributions:
  1. 1.
@@ 1. Introduction
    We find that, under idealized conditions of the Ethereum block building market, where builders have similar latency and orderflow access, builders are incentivized to compete against each other by bidding a significant portion of block value in the MEV-Boost auction, resulting in an equal market share (win rate). This fosters a decentralized and competitive Ethereum block building market.
  2. 2.
@@
    We identified which asymmetry contributes to centralization. Compared to the previous idealized scenario, our results indicate that the latency advantage alone within certain threshold does not significantly alter builders’ bidding incentives. However, latency asymmetry between builders leads to reduced proposer revenue (i.e., lower winning bid values) and decreased auction efficiency. These two findings contrast sharply with those in Wu et al. (2024b), suggesting that the construction of different strategy spaces can significantly impact builder incentives and auction outcomes.
  3. 3.
@@
    We demonstrate that advantage in private orderflow access significantly impacts builders’ bidding incentives. Builders with greater access to private orderflow are incentivized to bid a smaller portion of the block value while dominating the market and maintaining higher profit margins. Conversely, builders with limited orderflow access are forced to bid aggressively, using most of their block value for a marginal chance of winning. This disparity in orderflow access leads to centralization, oligopolistic behaviors, and auction inefficiencies in the Ethereum block building market.
2. Background and Related Works
@@
The Ethereum blockchain grows by one block every 12-second slot, as specified by its Proof-of-Stake consensus protocol Foundation (2022). At each slot, the protocol randomly selects a validator to act as the proposer, responsible for adding a new block to the blockchain at the start of that slot Ethereum (2022).
@@
MEV refers to the total value extractable from a blockchain state through transaction inclusion, exclusion, and ordering Daian et al. (2019), which are often carried out by MEV searchers under the current PBS framework with varying strategies Qin et al. (2022). The MEV supply chain outlines this process: Searchers share extracted MEV with builders by offering tips or payments for prioritized transaction inclusion, while builders share part of their MEV profits with proposers through bidding in the MEV-Boost auction. Vertical integration occurs when builders optimize this process by directly controlling searcher activity and relay infrastructure to gain an edge in MEV-Boost auctions with advantages in latency and access to searcher orderflow.
@@
Although Proposer-Builder Separation (PBS) has not yet been formally integrated into the Ethereum consensus protocol, MEV-Boost Flashbots (2022) was developed as an out-of-protocol solution to enhance decentralization. MEV-Boost introduces relays, trusted third-party intermediaries, to facilitate a fair exchange between the proposer and builders. Relays ensure that the proposer receives the payment from the winning builder, while builders are protected from having their MEV opportunities revealed and stolen. The proposer can either produce the block locally and extract MEV themselves, or use MEV-Boost to outsource block production and earn MEV revenue from builders. MEV-Boost auctions align with the 12-second slot interval: to get their block selected by the proposer of slot $n$, builders begin to submit their blocks along with bids to relays at the start of slot $n-1$. Relays validate these bids and make them available to the proposer, who typically terminates the auction by signing the block header with the highest bid and returning it the to corresponding relay, which then publishes the full block to the network.
@@
Apart from Wu et al. (2024b), previous research on builders’ behaviors in MEV-Boost auctions has explored various empirical, theoretical, and game-theoretic perspectives. Wahrstätter et al. (2023) conducted a longitudinal empirical study, revealing a positive correlation between bid timing and value in block production. Neuder (2023) identified strategic behaviors like bid erosion and bid shielding using empirical data, while Thiery (2023) confirmed diverse bidding patterns among builders. Öz et al. (2024) empirically highlighted key factors behind successful block wins, such as exclusive access to private orderflow providers. Yang et al. (2024) empirically confirmed that private orderflow serves as a significant market entry barrier for new builders, impacting efficiency and competitiveness in the market. Gupta et al. (2023) theoretically demonstrated that builders with high-value orderflow tend to perform better in MEV-Boost auctions and highlighted the centralizing effect of private orderflow in the Ethereum block building market. These findings were also empirically validated by Always (2023). Pai and Resnick (2023) further proved that latency advantages give builders a competitive edge. Wu et al. (2024a) introduced a game-theoretic model alongside exemplary bidding strategies showcasing the importance of both latency and orderflow access with simulations.
3. Model
The MEV-Boost auction was initially modelled in Wu et al. (2024a) and Wu et al. (2024b). In the current study, we enrich this model with the introduction of meta strategies and the construction of a meta game that better captures the practical aspects of the auction and the builders’ strategic decisions. To keep the paper self-contained, we first recap the main model’s components and parameters.
@@ 3. Model
Finally, the auction interval is defined as $[0,T]$, where $T$ denotes the time when the proposer selects the highest winning bid. Instead of $T$ being exactly equal to 12 seconds as expected, the winning bid is typically selected around $T=13$ seconds according to empirical data Wahrstätter (2023); Always (2024) due to factors such as the proposer’s latency or strategic behaviors Schwarz-Schilling et al. (2023). Thus, we draw $T$ randomly from a normal distribution with mean $13$ and standard deviation $\sigma$ [^2] We set $\sigma=0.1$ in simulations but the results are analogous across different values.
3.1. Model Calibration
@@
We calibrate the model against 75 days of mempool data Flashbots (2024b) and on-chain data Flashbots and Dune (2024) maintained by Flashbots and Dune Analytics covering the period from July 11 to September 23, 2024. Specifically, we consider $n=10$ players in the MEV-Boost auction game, as the top 10 builders build 97.51% of all blocks built via MEV-Boost. Each block contains approximately 105 public transactions, with the MEV of each transaction valued at approximately 0.00019 ETH, collectively representing nearly 31% of the total block value.
@@
As a result, we assume that private orderflow access among builders is randomly distributed within a domain that we estimate from on-chain data Flashbots and Dune (2024). This ensures that the volume of private orderflow in the winning block matches expectations from on-chain metrics. For the top 10 builders, private orderflow accounts for 10.92% (by builder blockbeelder) to 36.38% (by builder Titan) of transactions in their winning blocks. Accordingly, we infer that 36.38% represents the maximum possible private orderflow, corresponding to 100% access probability, while 10.92% marks the minimum, equating to a 30% probability. Thus, we model private orderflow access probabilities, $\theta_{i}$, as uniformly distributed between 30% and 100% (equivalently $[0.3,1]$).
3.2. Meta-game and meta-strategies
To prevent proposers from locally building blocks using public mempool transactions, builders are incentivized to bid at least the value of the public signal Wahrstätter (2024). Consequently, the builder’s strategic decision is how much of their private signal to include in their bid at time $t$. However, since players can bid any portion of their private signal, the strategy profile becomes infinite, making this game intractable.
[…]
@@ 3.2. Meta-game and meta-strategies
The above definitions assume that builders are not willing to bid in a way that leads to a negative profit. However, in current practice, builders sometimes subsidize their bids to win the auction to maintain a market share and expect better performance in the future Yang et al. (2024); Öz et al. (2024). We introduce this assumption here to simplify the analysis, focusing on non-repeated auction games.
@@
Table 1. Meta strategies and bid values $s_{i}(x_{i,t},m_{i}).$
@@ 4. Empirical Games
In this way, we can simplify the analysis by transforming the game into an anonymous game, where a player’s payoff depends only on the number of opponents adopting each strategy Wellman et al. (2024). This approach allows us to represent a strategy profile as a vector containing the count of players choosing each strategy, thereby reducing the number of profiles from $3^{10}$ to $\binom{10+\lvert S\rvert-1}{10}=66$.
We then replace the original payoff matrix with a heuristic payoff table (HPT) Walsh et al. (2002); Tuyls et al. (2018), where the payoffs of each meta strategy are stored as a function of the number of players using it. However, despite the players’ probabilities being drawn from the same prior distribution, their realized access may differ, leading to variability in payoffs even for identical strategy choices. This inherent asymmetry complicates the application of standard HPT methods, which assume symmetry across players Tuyls et al. (2018).
@@
Table 2. Empirical games.
@@
In the previous game setting, we assumed that all players had uniform latency. However, in practice, builders often experience different latencies Thiery (2023); Öz et al. (2024). In this section, we consider a scenario where players share the same distribution for private orderflow access but differ in their latencies. Specifically, we model two distinct latency groups: one group consisting of 5 players with low latency and another group consisting of 5 players with high latency.
This more realistic scenario allows us to model the game as a role-symmetric game Wellman et al. (2024), where players are categorized into roles based on their latency, and their payoffs depend on the strategies used within their role. In this case, there are two roles: $r_{l}$ for the low-latency group and $r_{h}$ for the high-latency group.
@@
The final variant considers differences in orderflow access probabilities among builders. In practice, private orderflow access is not uniform across all players; some builders have higher probabilities of accessing valuable orderflow than others Öz et al. (2024). In this variant, we model the game with 5 players having high private orderflow access probabilities and 5 players having low access probabilities, while all players have the same latency.
@@
4.1. $\alpha$-Rank
To solve the games discussed above, we utilize the $\alpha$-Rank algorithm Omidshafiei et al. (2019). The $\alpha$-Rank algorithm models a stochastic evolutionary process to capture the selection-mutation dynamics of a population’s strategy choices. This method provides a dynamic solution to the game by understanding the agents’ behaviors and predicting their convergence towards equilibrium. The solution (equilibrium), presented as a ranking of strategy profiles, corresponds to the stationary probabilities, also referred to as the mass, within the unique stationary distribution of the $\alpha$-Rank Markov Chain, which captures the evolutionary stable probabilities of the strategy profiles. We denote the stationary distribution as $\pi$ and the mass of a strategy profile $m^{k}$ as $\pi_{k}$. To convey the equilibrium clearly, we represent it as the average number of players using each strategy across all profiles, weighted by their respective mass. This representation effectively captures the expected frequency with which each strategy is employed by the players in the long run.

### v2

@@
Abstract
Block production on the Ethereum blockchain has adopted an auction-based mechanism known as Proposer-Builder Separation (PBS), where validators outsource block creation to builders competing in MEV-Boost auctions for Maximal Extractable Value (MEV) rewards. We employ empirical game-theoretic analysis based on simulations to examine how advantages in latency and MEV access shape builder strategic bidding and auction outcomes. We find that a small set of dominant builders leverage these advantages, consolidating power, reducing auction efficiency, and heightening centralization. Our results underscore the need for fair MEV distribution and sustained efforts to promote decentralization in Ethereum’s block building market.
1 Introduction
Modeling and analyzing strategic behavior in multi-agent systems is crucial for understanding how incentives drive agents’ decisions and shape overall system outcomes. Many applications rely on equilibrium analysis to predict how self-interested agents will act under various constraints. However, the realism of these predictions hinges on the breadth of agents’ strategy spaces: when strategies are constrained to be narrow, hand-crafted heuristics, interactions can be oversimplified, resulting in findings that fail to generalize or match real-world observations.
In this paper, we adopt the meta game framework [29] to capture a broader set of agent behaviors, defined by meta strategies [24], and thus, model multi-agent decision-making more comprehensively. This approach allows agents to explore a wider range of strategies, uncovering behaviors that might remain hidden when relying solely on pre-defined or empirically derived heuristics. We demonstrate the effectiveness of this framework in the context of Decentralized Finance (DeFi), specifically the mechanism Proposer-Builder Separation (PBS) and the Ethereum block building market [1, 8]—while emphasizing that our methodology is broadly applicable to other multi-agent environments.
DeFi provides a rich case study. In Ethereum’s current block production process, PBS decouples the block proposal and block building processes, allowing the proposer—a validator selected by the protocol to add a new block—to outsource the task of block building and MEV extraction to specialized entities known as builders [8]. Profit-driven builders compete in English-style auctions, known as MEV-Boost auctions [9], for the right to produce new blocks and earn potential Maximal Extractable Value (MEV) rewards [7]. Ideally, each builder would have equal network access and resources, creating a fair, competitive market with an efficient auction. In practice, however, builders seek to secure advantages—such as exclusive private transactions (orderflow) and reduced latency—to gain a competitive edge. Consequently, as of April 2025, two builders (beaverbuild and Titan) dominated over 90% of the auctions [13, 30], reaping high profits and raising centralization concerns.
Despite empirical studies revealed such oligopoly in the Ethereum block building market [38, 22], insights into how latency and private orderflow asymmetries shape bidding incentives in MEV-Boost auctions remain limited. Prior work restricts builders to empirically validated or narrowly adaptive heuristic strategies [36, 35], yielding limited insight into emerging behaviors and incentives. Counterintuitively, [35] finds that builders exhibit mild collusion that leads to price-fixing, even with private orderflow asymmetries.
In this paper, we address this gap by considering MEV-Boost auctions within the meta-game framework that broadens the strategy space, capturing more comprehensive bidding behaviors (Section 3). We investigate builders’ strategic bidding incentives under various scenarios (Section 4) and show that advantages in latency and private orderflow can significantly shift bidding incentives and auction outcomes. These results align more closely with real-world observations and offer novel insights into the key dynamics and builders’ incentives that drive the outcomes in current MEV-Boost auctions (Section 5). In particular, our paper makes the following contributions.
@@ 1 Introduction
First, we find that, under idealized conditions, where builders have similar latency and orderflow access, builders are incentivized to compete against each other by bidding a significant portion of block value, resulting in an equal market share (win rate). This fosters a decentralized and competitive Ethereum block building market.
@@
Second, we identify which asymmetry contributes to centralization. Compared to the previous idealized scenario, our results indicate that the latency advantage alone within certain threshold does not significantly alter builders’ bidding incentives or contribute to centralization. However, latency asymmetry can lead to reduced proposer revenue (i.e., lower winning bid values) and decreased auction efficiency. These two findings contrast sharply with those in [35], highlighting the importance of strategy space design.
@@
Third, we demonstrate that advantage in private orderflow access significantly impacts builders’ bidding incentives. Builders with greater access to private orderflow are incentivized to bid a smaller portion of the block value while dominating the market and maintaining higher profit margins. Conversely, builders with limited orderflow access are forced to bid aggressively. This disparity in orderflow access leads to centralization, oligopolistic behaviors, and auction inefficiencies in the Ethereum block building market.
Finally, we conclude our analysis by discussing the broader challenges in the Ethereum block building market and the implications of our approach and analysis for future studies (Section 6).
2 Background and Related Works
@@
The Ethereum blockchain grows by one block every 12-second slot, as specified by its Proof-of-Stake consensus protocol [14]. At each slot, the protocol randomly selects a validator to act as the proposer, responsible for adding a new block to the blockchain at the start of that slot [6].
@@
MEV refers to the total value extractable from a blockchain state through transaction inclusion, exclusion, and ordering [3], which are often carried out by MEV searchers under the current PBS framework with varying strategies [26, 25, 20, 37]. The MEV supply chain outlines this process: Searchers share extracted MEV with builders by offering tips or payments for prioritized transaction inclusion, while builders share part of their MEV profits with proposers through bidding in the MEV-Boost auction. Vertical integration occurs when builders optimize this process by directly controlling searcher activity and relay infrastructure to gain an edge in MEV-Boost auctions with advantages in latency and access to searcher orderflow.
@@
Although Proposer-Builder Separation (PBS) has not yet been formally integrated into the Ethereum consensus protocol, MEV-Boost [9] was developed as an out-of-protocol solution to enhance decentralization. MEV-Boost introduces relays, trusted third-party intermediaries, to facilitate a fair exchange between the proposer and builders. Relays ensure that the proposer receives the payment from the winning builder, while builders are protected from having their MEV opportunities revealed and stolen. The proposer can either produce the block locally and extract MEV themselves, or use MEV-Boost to outsource block production and earn MEV revenue from builders. MEV-Boost auctions align with the 12-second slot interval: to get their block selected by the proposer of slot $n$, builders begin to submit their blocks along with bids to relays at the start of slot n1n-1. Relays validate these bids and make them available to the proposer, who typically terminates the auction by signing the block header with the highest bid and returning it the to corresponding relay, which then publishes the full block to the network.
@@
Apart from [35], previous research on builders’ behaviors in MEV-Boost auctions has explored various empirical, theoretical, and game-theoretic perspectives. [32, 16] conducted a longitudinal empirical study, revealing a positive correlation between bid timing and value in block production and highlighting centralization concerns in the Ethereum builder market. [19] identified strategic behaviors like bid erosion and bid shielding using empirical data, while [28] confirmed diverse bidding patterns among builders. [22] empirically highlighted key factors behind successful block wins, such as exclusive access to private orderflow providers. [38] empirically confirmed that private orderflow serves as a significant market entry barrier for new builders, impacting efficiency and competitiveness in the market. [15] theoretically demonstrated that builders with high-value orderflow tend to perform better in MEV-Boost auctions and highlighted the centralizing effect of private orderflow in the Ethereum block building market. These findings were also empirically validated by [4]. [23] further proved that latency advantages give builders a competitive edge. [36] introduced a game-theoretic model alongside exemplary bidding strategies showcasing the importance of both latency and orderflow access with simulations.
3 Model
The MEV-Boost auction was initially modelled in [36] and [35]. In the current study, we enrich this model with the introduction of meta strategies and the construction of a meta game that better captures the practical aspects of the auction and the builders’ strategic decisions. To keep the paper self-contained, we first recap the main model’s components and parameters.
@@ 3 Model
Finally, the auction interval is defined as $[0,T]$, where $T$ denotes the time when the proposer selects the highest winning bid. Instead of $T$ being exactly equal to 12 seconds as expected, the winning bid is typically selected around $T=13$ seconds according to empirical data [30, 5] due to factors such as the proposer’s latency or strategic behaviors [27]. Thus, we draw $T$ randomly from a normal distribution with mean $13$ and standard deviation $\sigma$. [^1] We set $\sigma=0.1$ in simulations but the results are analogous across different values.
3.1 Model Calibration
@@
We calibrate the model against 75 days of mempool data [11] and on-chain data [13] maintained by Flashbots and Dune Analytics covering the period from July 11 to September 23, 2024. Specifically, we consider $n=10$ players in the MEV-Boost auction game, as the top 10 builders build 97.51% of all blocks built via MEV-Boost. Each block contains approximately 105 public transactions, with the MEV of each transaction valued at approximately 0.00019 ETH, collectively representing nearly 31% of the total block value.
@@
As a result, we assume that private orderflow access among builders is randomly distributed within a domain that we estimate from on-chain data [13]. This ensures that the volume of private orderflow in the winning block matches expectations from on-chain metrics. For the top 10 builders, private orderflow accounts for 10.92% (by builder blockbeelder) to 36.38% (by builder Titan) of transactions in their winning blocks. Accordingly, we infer that 36.38% represents the maximum possible private orderflow, corresponding to 100% access probability, while 10.92% marks the minimum, equating to a 30% probability. Thus, we model private orderflow access probabilities, $\theta_{i}$, as uniformly distributed between 30% and 100% (equivalently $[0.3,1]$).
3.2 Meta-game and meta-strategies
To prevent proposers from locally building blocks using public mempool transactions, builders are incentivized to bid at least the value of the public signal [31]. Consequently, the builder’s strategic decision is how much of their private signal to include in their bid at time $t$. However, since players can bid any portion of their private signal, the strategy profile becomes infinite, making this game intractable.
To address this complexity, we introduce a meta game [29], which simplifies the original game by focusing only on a set of meta strategies that summarise all possible strategies [24]. In the context of MEV-Boost auctions, we define the meta strategy set as $S=\{\emph{"conservative", "moderate", "aggressive"}\}$, or $S=\{m_{1},m_{2},m_{3}\}$. In particular, we divide the $[0,1]$ interval, representing the fraction of private signal included in a bid, into three sub-intervals that correspond to these strategies. Thus, the bid of a builder, $i$, is given by (cf. Table˜1)
@@ 3.2 Meta-game and meta-strategies
The above definitions assume that builders are not willing to bid in a way that leads to a negative profit. However, in current practice, builders sometimes subsidize their bids to win the auction to maintain a market share and expect better performance in the future [38, 22]. We introduce this assumption here to simplify the analysis, focusing on non-repeated auction games.
@@
Table 1: Meta strategies and bid values $s_{i}(x_{i,t},m_{i}).$
@@ 4 Game Settings
In this way, we can simplify the analysis by transforming the game into an anonymous game, where a player’s payoff depends only on the number of opponents adopting each strategy [34]. This approach allows us to represent a strategy profile as a vector containing the count of players choosing each strategy, thereby reducing the number of profiles from $3^{10}$ to $\binom{10+\lvert S\rvert-1}{10}=66$.
We then replace the original payoff matrix with a heuristic payoff table (HPT) [33, 29], where the payoffs of each meta strategy are stored as a function of the number of players using it. However, despite the players’ probabilities being drawn from the same prior distribution, their realized access may differ, leading to variability in payoffs even for identical strategy choices. This inherent asymmetry complicates the application of standard HPT methods, which assume symmetry across players [29].
@@
Table 2: Game settings.
@@
In the previous game setting, we assumed that all players had uniform latency. However, in practice, builders often experience different latencies [28, 22]. In this section, we consider a scenario where players share the same distribution for private orderflow access but differ in their latencies. Specifically, we model two distinct latency groups: one group consisting of 5 players with low latency and another group consisting of 5 players with high latency.
This more realistic scenario allows us to model the game as a role-symmetric game [34], where players are categorized into roles based on their latency, and their payoffs depend on the strategies used within their role. In this case, there are two roles: $r_{l}$ for the low-latency group and $r_{h}$ for the high-latency group.
@@
The final variant considers differences in orderflow access probabilities among builders. In practice, private orderflow access is not uniform across all players; some builders have higher probabilities of accessing valuable orderflow than others [22]. In this variant, we model the game with 5 players having high private orderflow access probabilities and 5 players having low access probabilities, while all players have the same latency.
@@
4.1 $\alpha$-Rank
To solve the games discussed above, we utilize the $\alpha$-Rank algorithm [21]. The $\alpha$-Rank algorithm models a stochastic evolutionary process to capture the selection-mutation dynamics of a population’s strategy choices. This method provides a dynamic solution to the game by understanding the agents’ behaviors and predicting their convergence towards equilibrium. The solution (equilibrium), presented as a ranking of strategy profiles, corresponds to the stationary probabilities, also referred to as the mass, within the unique stationary distribution of the $\alpha$-Rank Markov Chain, which captures the evolutionary stable probabilities of the strategy profiles. We denote the stationary distribution as $\pi$ and the mass of a strategy profile $m^{k}$ as $\pi_{k}$. To convey the equilibrium clearly, we represent it as the average number of players using each strategy across all profiles, weighted by their respective mass. This representation effectively captures the expected frequency with which each strategy is employed by the players in the long run.